Calculate the normal and the geodesic curvatures of the following curves on
the given surfaces:
(a) The circle γ(t) = (cost, sin t, 1) on the paraboloid σ(u, v) = (u, v, u^2 + v^2).
1
Expert's answer
2020-12-02T18:32:01-0500

The normal curvature, kn, is the curvature of the curve projected onto the plane containing the curve's tangent T and the surface normal u; the geodesic curvature, kg, is the curvature of the curve projected onto the surface's tangent plane

and normal curvature, "K_n"


The unit normal vector is given by (http://mathonline.wikidot.com/the-frenet-serret-formulas): "N(s)=\frac{T'(s)}{||T'(s)||}" , where "T(s)=\frac{\vec{r}'(s)}{||\vec{r}'(s)||}"

where "r(s)" is an arc-length parametrization of "\vec{r}(t)"

"\vec{r}(t)".

(a) We point out that one has an arc-length parametrization in this case. The length of the circle is "2\pi" and it corresponds to the interval "[0,2\pi)" for possible values of "t". Thus, t=s. We have:

"r'(t)=(-sin(t),cos(t),0)\Rightarrow ||\vec{r}'(t)||=1."

"T(t)=(-sin(t),cos(t),0)"; "T'(t)=(-cos(t),-sin(t),0)";

"N(t)=(-cos(t),sin(t),0)".

The geodesic curvature is (https://encyclopediaofmath.org/wiki/Geodesic_curvature):

"k_g=\frac{(r',r'',n)}{|r'|^3}" , where "n" is a normal vector for the surface. The normal vector is:

"n=(2u,2v,-1)" (see https://mathworld.wolfram.com/NormalVector.html).

After substitution of the curve equation, we get that "n=(2cos\,t,2sin\,t,-1)" on the points of the curve. Thus,

"k_g=\begin{vmatrix}
 -sin\,t & cos\,t &0 \\
 -cos\,t & -sin\,t &0\\
2cos\,t & 2sin\,t &-1
\end{vmatrix}=-1"


The normal curvature is related to the second fundamental form, and an expression for it is


"k_n=L \dot{u}^2+2M\dot{u}\dot{v}+N\dot{v}^2"

Where there is the surface on the paraboloid "\sigma (u,v)=(u, v, u^2+v^2)" then

"L=\sigma_{uu} \cdot \vec{N},\ M=\sigma_{uv} \cdot \vec{N}, L=\sigma_{vv} \cdot \vec{N}," where "\vec{N}" is the normal vector to the surface.

But on the paraboloid, "z=0". So, "\sigma_{uu}=\sigma_{uv}=\sigma_{vv}=0" and also "L=M=N=0" thus "k_n=0"

Hence the solution for "K_g" and "K_n"